# A Geometric View of Data Complexity: Efficient Local Intrinsic Dimension Estimation with Diffusion Models

### v1

@@ Abstract
High-dimensional data commonly lies on low-dimensional submanifolds, and estimating the local intrinsic dimension (LID) of a datum – i.e. the dimension of the submanifold it belongs to – is a longstanding problem. LID can be understood as the number of local factors of variation: the more factors of variation a datum has, the more complex it tends to be. Estimating this quantity has proven useful in contexts ranging from generalization in neural networks to detection of out-of-distribution data, adversarial examples, and AI-generated text. The recent successes of deep generative models present an opportunity to leverage them for LID estimation, but current methods based on generative models produce inaccurate estimates, require more than a single pre-trained model, are computationally intensive, or do not exploit the best available deep generative models, i.e. diffusion models (DMs). In this work, we show that the Fokker-Planck equation associated with a DM can provide a LID estimator which addresses all the aforementioned deficiencies. Our estimator, called FLIPD, is compatible with all popular DMs, and outperforms existing baselines on LID estimation benchmarks. We also apply FLIPD on natural images where the true LID is unknown. Compared to competing estimators, FLIPD exhibits a higher correlation with non-LID measures of complexity, better matches a qualitative assessment of complexity, and is the only estimator to remain tractable with high-resolution images at the scale of Stable Diffusion.
@@ 1 Introduction
Figure 1: (Left) A cartoon illustration showing that LID is a natural measure of relative complexity. We depict two manifolds of MNIST digits, corresponding to 1s and 8s, as 1-dimensional and 2-dimensional submanifolds of $\mathbb{R}^{3}$, respectively. The relatively simpler manifold of 1s exhibits a single factor of variation (“tilt”), whereas 8s have an additional factor of variation (“disproportionality”). (Right) The 4 lowest- and highest-LID datapoints from a subsample of LAION Aesthetics, as measured by our method, FLIPD, applied to Stable Diffusion v1.5. FLIPD scales efficiently to large models on high-dimensional data, and aligns closely with subjective complexity.
@@
We address all these issues by showing how LID can be efficiently estimated using only a single pre-trained diffusion model (DM) by building on LIDL [60], a model-based estimator. LIDL operates by convolving data with different levels of Gaussian noise, training a normalizing flow [48, 16, 17] for each level, and fitting a linear regression using the log standard deviation of the noise as a covariate and the corresponding log density (of the convolution) evaluated at $x$ as the response; the resulting slope is an estimate of $\text{LID}(x)-D$, thanks to a surprising result linking Gaussian convolutions and LID. We first show how to adapt LIDL to DMs in such a way that only a single DM is required (rather than many normalizing flows). Directly applying this insight leads to LIDL estimates that require one DM but several calls to an ordinary differential equation (ODE) solver; we also show how to circumvent this with an alternative ODE that computes all the required log densities in a single solve. We then argue that the slope of the regression in LIDL aims to capture a rate of change which, for DMs, can be evaluated directly thanks to the Fokker-Planck equation. The resulting estimator, which we call FLIPD, [^1] Pronounced as “flipped”, the acronym is a rearrangement of “FP” from Fokker-Planck and “LID”. is highly efficient and circumvents the need for an ODE solver.
Our contributions are: $(i)$ showing how DMs can be efficiently combined with LIDL in a way which requires a single call to an ODE solver; $(ii)$ leveraging the Fokker-Planck equation to propose FLIPD, thus improving upon the estimator and circumventing the need for an ODE solver altogether; $(iii)$ motivating FLIPD theoretically; $(iv)$ introducing a suite of LID estimation benchmark tasks highlighting that the previously reported good performance of other estimators is due in part to the simplicity of the tasks over which they were evaluated; $(v)$ demonstrating that FLIPD tends to outperform existing baselines, particularly in high-dimensional settings, while being much more computationally efficient – FLIPD is the first LID estimator to scale to high-resolution images ($\sim 10^{6}$ dimensions), and can be used with Stable Diffusion; and $(vi)$ showing that when applied to natural images, compared to competing alternatives, FLIPD is more aligned both with other measures of complexity such as PNG compression length, and with qualitative assessments of complexity.
@@ 2.1 Diffusion Models
where $\hat{p}(\cdot,1)$ can be evaluated since it is a known Gaussian, and where $v(x,t)\coloneqq f(x,t)-g^{2}(t)\hat{s}(x,t)/2$.
@@
Note that the cost of computing $\nabla v(\hat{x}_{t},t)$ for a particular $\hat{x}_{t}$ amounts to $\mathcal{O}(D)$ function evaluations of $\hat{s}$ (since $D$ calls to a Jacobian-vector-product routine are needed [6]). Although this is not prohibitively expensive for a single $\hat{x}_{t}$, in order to compute the integral in Equation 5 in practice, $(\hat{x}_{t})_{t\in[t_{0},1]}$ must be discretized into a trajectory of length $N$. Deterministic density evaluation is thus computationally intractable as it costs $\mathcal{O}(ND)$ function evaluations. The Hutchinson trace estimator [28] – which states that for $M\in\mathbb{R}^{D\times D}$, $\operatorname{tr}(M)=\mathbb{E}_{\varepsilon}[\varepsilon^{\top}M\varepsilon]$, where $\varepsilon\in\mathbb{R}^{D}$ has mean 00 and covariance $I_{D}$ – is thus commonly used for stochastic density estimation; approximating the expectation with $k$ samples from $\varepsilon$ results in a cost of $\mathcal{O}(Nk)$, which is much faster than deterministic density evaluation when $k\ll D$.
@@ 2.2 Local Intrinsic Dimension and How to Estimate It
Various definitions of intrinsic dimension exist [27, 20, 36], here we follow the standard one from geometry: a $d$-dimensional manifold is a set which is locally homeomorphic to $\mathbb{R}^{d}$. For a given disjoint union of manifolds and a point $x$ in this union, the local intrinsic dimension of $x$ is the dimension of the manifold it belongs to. Note that LID is not an intrinsic property of the point $x$, but rather a property of $x$ with respect to the manifold that contains it. Intuitively, $\text{LID}(x)$ corresponds to the number of factors of variation present in the manifold containing $x$, and it is thus a natural measure of the relative complexity of $x$, as illustrated in Figure 1.
@@
for some constant $c$. If we could evaluate $\log\varrho(x,\delta)$ for various values of $\delta$, this would provide an avenue for estimating $\text{LID}(x)$: set some values $\delta_{1},\dots,\delta_{m}$, fit a linear regression using $\{(\delta_{i},\log\varrho(x,\delta_{i}))\}_{i=1}^{m}$ (with $\delta$ as the covariate and $\log\varrho(x,\delta)$ as the response), and let $\hat{\beta}_{x}$ be the corresponding slope. It follows that $\hat{\beta}_{x}$ estimates $\text{LID}(x)-D$, so that $\text{LID}(x)\approx D+\hat{\beta}_{x}$ is a sensible estimator of local intrinsic dimension.
@@
The only works we are aware of that leverage DMs for LID estimation are those of Stanczuk et al. [57], and Horvat and Pfister [25]. The latter modifies the training procedure of DMs, so we focus on the former since we see compatibility with existing pre-trained models as an important requirement for DM-based LID estimators. Stanczuk et al. consider variance-exploding DMs, where $f=0$. They show that, as $t\searrow 0$, the score function $s(x,t)$ points orthogonally towards the manifold containing $x$, or more formally, it lies in the normal space of this manifold at $x$. They thus propose the following LID estimator which we refer to as the normal bundle (NB) estimator: first run Equation 1 until time $t_{0}$ starting from $x$, and evaluate $\hat{s}(\cdot,t_{0})$ at the resulting value; then repeat this process $K$ times and stack the $K$ resulting $D$-dimensional vectors into a matrix $S(x)\in\mathbb{R}^{D\times K}$. The idea here is that if $t_{0}$ is small enough and $K$ is large enough, the columns of $S(x)$ span the normal space of the manifold at $x$, suggesting that the rank of this matrix estimates the dimension of this normal space, namely $D-\text{LID}(x)$. Finally, they estimate $\text{LID}(x)$ as:
@@
Although this estimator addresses some of the limitations of LIDL, it remains computationally expensive – computing $S(x)$ requires $K$ function evaluations with the authors recommending $K=4D$ – and as we will see later, it does not always produce reliable estimates.
@@ 3 Method
Although Tempczyk et al. [60] only used normalizing flows in LIDL, they did point out that these models could be swapped for any other generative model admitting density evaluation. Indeed, one could trivially train $m$ DMs and replace the flows with them. Throughout this section we will assume access to a pre-trained DM such that $f(x,t)=b(t)x$ for a function $b:[0,1]\rightarrow\mathbb{R}$. This choice implies that the transition kernel $p_{t|0}$ associated with Equation 1 is Gaussian [51]:
@@
where $\psi,\sigma:[0,1]\rightarrow\mathbb{R}$. We also assume that $b$ and $g$ are such that $\psi$ and $\sigma$ are differentiable and such that $\lambda(t)\coloneqq\sigma(t)/\psi(t)$ is injective. This setting encompasses all DMs commonly used in practice, including variance-exploding, variance-preserving (of which DDPM [24] is a discretized instance), and sub-variance-preserving [56]. In Appendix A we include explicit formulas for $\psi(t)$, $\sigma^{2}(t)$, and $\lambda(t)$ for these particular DMs.
@@ 3.1 LIDL with a Single Diffusion Model
As opposed to the several normalizing flows used in LIDL which are individually trained on datasets with different levels of noise added, a single DM already works by convolving data with various noise levels and allows density evaluation of the resulting noisy distributions (Equation 5). Hence, we make the observation that LIDL can be used with a single DM. All we need is to relate $\varrho(x,\delta)$ to the density of the DM for any given $x$ and $\delta$. In the case of variance-exploding DMs, $\psi(t)=1$, so we can easily use the defining property of the transition kernel in Equation 10 to get
@@
which equals $\varrho(x,\delta)$ from Equation 6 when we choose $t=\sigma^{-1}(e^{\delta})$. [^3] Note that we treat positive and negative superindices differently: e.g. $\sigma^{-1}$ denotes the inverse function of $\sigma$, not $1/\sigma$; on the other hand $\sigma^{2}$ denotes the square of $\sigma$, not $\sigma\circ\sigma$. In turn, we can use LIDL with a single variance-exploding DM by evaluating each $\hat{p}(x,\sigma^{-1}(e^{\delta_{i}}))$ through Equation 5. This idea extends beyond variance-exploding DMs; in Appendix B.1 we show that for any arbitrary DM with transition kernel as in Equation 10, it holds that
@@
where $t(\delta)\coloneqq\lambda^{-1}(e^{\delta})$, thus showing that LIDL can be used with a single DM.
3.2 An Efficient Implementation of LIDL with Diffusion Models
@@
Using Equation 12 with LIDL still involves computing $\log\hat{p}\big{(}\psi(t(\delta_{i}))x,t(\delta_{i})\big{)}$ through Equation 5 for each $i=1,\dots,m$ before running the regression. Since each of the corresponding ODEs in Equation 4 starts at a different time $t_{0}=t(\delta_{i})$ and is evaluated at a different point $\psi(t(\delta_{i}))x$, this means that a different ODE solver call would have to be used for each $i$, resulting in a prohibitively expensive procedure. By leveraging the Fokker-Planck equation associated with Equation 1, we show in Appendix B.2 that, for DMs with transition kernel as in Equation 10,
@@
Assuming without loss of generality that $\delta_{1}<\dots<\delta_{m}$, solving the ODE
${\textnormal{d}}\log\hat{\varrho}(x,\delta)=\nu\big{(}t(\delta);\hat{s},x\big{)}{\textnormal{d}}\delta,\quad\log\hat{\varrho}(x,\delta_{1})=0,$ (14)
@@
from $\delta_{1}$ to $\delta_{m}$ produces the trajectory $(\log\hat{\varrho}(x,\delta))_{\delta\in[\delta_{1},\delta_{m}]}$. Since $\nu(t(\delta);\hat{s},x)$ does not depend on $\hat{\varrho}(x,\delta)$, when $\hat{s}=s$ the solution will be such that $\log\hat{\varrho}(x,\delta)=\log\varrho(x,\delta)+c_{\text{init}}$ for some constant $c_{\text{init}}$ that depends on the initial condition but not on $\delta$. Thus, while setting the initial condition to 00 might at first appear odd, we can fit a regression as in LIDL, using $\{(\delta_{i},\log\hat{\varrho}(x,\delta_{i}))\}_{i=1}^{m}$, and $c_{\text{init}}$ will be absorbed in the intercept without affecting the slope. In other words, the initial condition is irrelevant, and we can use LIDL with DMs by calling a single ODE solver on Equation 14.
3.3 FLIPD: A Fokker-Planck-Based LID Estimator

### v2

@@ Abstract
High-dimensional data commonly lies on low-dimensional submanifolds, and estimating the local intrinsic dimension (LID) of a datum – i.e. the dimension of the submanifold it belongs to – is a longstanding problem. LID can be understood as the number of local factors of variation: the more factors of variation a datum has, the more complex it tends to be. Estimating this quantity has proven useful in contexts ranging from generalization in neural networks to detection of out-of-distribution data, adversarial examples, and AI-generated text. The recent successes of deep generative models present an opportunity to leverage them for LID estimation, but current methods based on generative models produce inaccurate estimates, require more than a single pre-trained model, are computationally intensive, or do not exploit the best available deep generative models: diffusion models (DMs). In this work, we show that the Fokker-Planck equation associated with a DM can provide an LID estimator which addresses the aforementioned deficiencies. Our estimator, called FLIPD, is easy to implement and compatible with all popular DMs. Applying FLIPD to synthetic LID estimation benchmarks, we find that DMs implemented as fully-connected networks are highly effective LID estimators that outperform existing baselines. We also apply FLIPD to natural images where the true LID is unknown. Despite being sensitive to the choice of network architecture, FLIPD estimates remain a useful measure of relative complexity; compared to competing estimators, FLIPD exhibits a consistently higher correlation with image PNG compression rate and better aligns with qualitative assessments of complexity. Notably, FLIPD is orders of magnitude faster than other LID estimators, and the first to be tractable at the scale of Stable Diffusion.
@@ 1 Introduction
Figure 1: (Left) A cartoon illustration showing that LID is a natural measure of relative complexity. We depict two manifolds of MNIST digits, corresponding to 1s and 8s, as 1-dimensional and 2-dimensional submanifolds of $\mathbb{R}^{3}$, respectively. The relatively simpler manifold of 1s exhibits a single factor of variation (“tilt”), whereas 8s have an additional factor of variation (“disproportionality”). (Right) The 4 lowest- and highest-LID datapoints from a subsample of LAION-Aesthetics, as measured by our method, FLIPD, applied to Stable Diffusion v1.5. FLIPD scales efficiently to large models on high-dimensional data, and aligns closely with subjective complexity.
@@
We address these issues by showing how LID can be efficiently estimated using only a single pre-trained diffusion model (DM) by building on LIDL [61], a model-based estimator. LIDL operates by convolving data with different levels of Gaussian noise, training a normalizing flow [49, 17, 18] for each level, and fitting a linear regression using the log standard deviation of the noise as a covariate and the corresponding log density (of the convolution) evaluated at $x$ as the response; the resulting slope is an estimate of $\text{LID}(x)-D$, thanks to a surprising result linking Gaussian convolutions and LID. We first show how to adapt LIDL to DMs in such a way that only a single DM is required (rather than many normalizing flows). Directly applying this insight leads to LIDL estimates that require one DM but several calls to an ordinary differential equation (ODE) solver; we further show how to circumvent this with an alternative ODE that computes all the required log densities in a single solve. We then argue that the slope of the regression in LIDL aims to capture the rate of change of the marginal log probabilities in the diffusion process, which can be evaluated directly thanks to the Fokker-Planck equation. The resulting estimator, which we call FLIPD, [^1] Pronounced as “flipped”, the acronym is a rearrangement of “FP” from Fokker-Planck and “LID”. is highly efficient, and circumvents the need for an ODE solver. Notably, FLIPD is differentiable; this property opens up exciting avenues for future research as it enables backpropagating through LID estimates.
Our contributions are: $(i)$ showing how DMs can be efficiently combined with LIDL in a way which requires a single call to an ODE solver; $(ii)$ leveraging the Fokker-Planck equation to propose FLIPD, thus improving upon the estimator and circumventing the need for an ODE solver altogether; $(iii)$ motivating FLIPD theoretically; $(iv)$ introducing an expanded suite of LID estimation benchmark tasks that reveals gaps in prior evaluations, and specifically that other estimators do not remain accurate as the complexity of the manifold increases; $(v)$ demonstrating that when using fully-connected architectures for diffusion models, FLIPD outperforms existing baselines – especially as dimension increases – while being much more computationally efficient; $(vi)$ showing that when applied to natural images, despite varying across network architecture (i.e., fully-connected network or UNet [51, 67]), FLIPD estimates consistently align with other measures of complexity such as PNG compression length, and with qualitative assessments of complexity, highlighting that the LID estimates provided by FLIPD remain valid measures of relative image complexity; and $(vii)$ demonstrating that when applied to the latent space of Stable Diffusion, FLIPD can estimate LID for extremely high-resolution images ($\sim 10^{6}$ pixel dimensions) for the first time.
@@ 2.1 Diffusion Models
where $\hat{p}(\cdot,1)$ can be evaluated since it is a Gaussian, and where $v(x,t)\coloneqq f(x,t)-g^{2}(t)\hat{s}(x,t)/2$.
@@
Note that the cost of computing $\nabla v(\hat{x}_{t},t)$ for a particular $\hat{x}_{t}$ amounts to $\Theta(D)$ function evaluations of $\hat{s}$ (since $D$ calls to a Jacobian-vector-product routine are needed [6]). Although this is not prohibitively expensive for a single $\hat{x}_{t}$, in order to compute the integral in Equation 5 in practice, $(\hat{x}_{t})_{t\in[t_{0},1]}$ must be discretized into a trajectory of length $N$. If we denote by $F$ the cost of evaluating $v(\hat{x}_{t},t)$ – or equivalently, $\hat{s}(\hat{x}_{t},t)$ – deterministic density evaluation is $\Theta(NDF)$, which is computationally prohibitive. The Hutchinson trace estimator [29] – which states that for $M\in\mathbb{R}^{D\times D}$, $\operatorname{tr}(M)=\mathbb{E}_{\varepsilon}[\varepsilon^{\top}M\varepsilon]$, where $\varepsilon\in\mathbb{R}^{D}$ has mean 00 and covariance $I_{D}$ – is thus commonly used for stochastic density estimation; approximating the expectation with $k$ samples from $\varepsilon$ results in a cost of $\Theta(NkF)$, which is much faster than deterministic density evaluation when $k\ll D$.
@@ 2.2 Local Intrinsic Dimension and How to Estimate It
Various definitions of intrinsic dimension exist [28, 21, 37, 11]. Here we follow the standard one from geometry: a $d$-dimensional manifold is a set which is locally homeomorphic to $\mathbb{R}^{d}$. For a given disjoint union of manifolds and a point $x$ in this union, the local intrinsic dimension of $x$ is the dimension of the submanifold it belongs to. Note that LID is not an intrinsic property of the point $x$, but rather a property of $x$ with respect to the manifold that contains it. Intuitively, $\text{LID}(x)$ corresponds to the number of factors of variation present in the manifold containing $x$, and it is thus a natural measure of the relative complexity of $x$, as illustrated in Figure 1.
@@
for some constant $c$. If we could evaluate $\log\varrho(x,\delta)$ for various values of $\delta$, this would provide an avenue for estimating $\text{LID}(x)$: set some values $\delta_{1},\dots,\delta_{m}$, fit a linear regression using $\{(\delta_{i},\log\varrho(x,\delta_{i}))\}_{i=1}^{m}$ with $\delta$ as the covariate and $\log\varrho(x,\delta)$ as the response, and let $\hat{\beta}_{x}$ be the corresponding slope. It follows that $\hat{\beta}_{x}$ estimates $\text{LID}(x)-D$, so that $\text{LID}(x)\approx D+\hat{\beta}_{x}$ is a sensible estimator of local intrinsic dimension.
@@
The only works we are aware of that leverage DMs for LID estimation are those of Stanczuk et al. [58], and Horvat and Pfister [26]. The latter modifies the training procedure of DMs, so we focus on the former since we see compatibility with existing pre-trained models as an important requirement for DM-based LID estimators. Stanczuk et al. [58] consider variance-exploding DMs, where $f=0$. They show that, as $t\searrow 0$, the score function $s(x,t)$ points orthogonally towards the manifold containing $x$, or more formally, it lies in the normal space of this manifold at $x$. They thus propose the following LID estimator, which we refer to as the normal bundle (NB) estimator: first run Equation 1 until time $t_{0}$ starting from $x$, and evaluate $\hat{s}(\cdot,t_{0})$ at the resulting value; then repeat this process $K$ times and stack the $K$ resulting $D$-dimensional vectors into a matrix $S(x)\in\mathbb{R}^{D\times K}$. The idea here is that if $t_{0}$ is small enough and $K$ is large enough, the columns of $S(x)$ span the normal space of the manifold at $x$, suggesting that the rank of this matrix estimates the dimension of this normal space, namely $D-\text{LID}(x)$. Finally, they estimate $\text{LID}(x)$ as:
@@
Numerically, the rank is computed by performing a singular value decomposition (SVD) of $S(x)$, setting a threshold, and counting the number of singular values exceeding the threshold. Computing $S(x)$ requires $K$ function evaluations, and intuitively $K$ should be large enough to ensure the columns of $S(x)$ span the normal space at $x$; the authors thus propose using $K=4D$, and recommend always at least ensuring that $K>D$. Computing the NB estimator costs $\Theta(KF+D^{2}K)$, where $F$ again denotes the cost for evaluating $\hat{s}$. Thus, although the NB estimator addresses some of the limitations of LIDL, it remains computationally expensive in high dimensions.
@@ 3 Method
Although Tempczyk et al. [61] only used normalizing flows in LIDL, they did point out that these models could be swapped for any other generative model admitting density evaluation. Indeed, one could trivially train $m$ DMs and replace the flows with them. Throughout this section we provide a sequence of progressive improvements to this naïve application of LIDL with DMs, culminating with FLIPD. We assume access to a pre-trained DM such that $f(x,t)=b(t)x$ for a function $b:[0,1]\rightarrow\mathbb{R}$. This choice implies that the transition kernel $p_{t|0}$ associated with Equation 1 is Gaussian [52]:
@@
where $\psi,\sigma:[0,1]\rightarrow\mathbb{R}$. We also assume that $b$ and $g$ are such that $\psi$ and $\sigma$ are differentiable and such that $\lambda(t)\coloneqq\sigma(t)/\psi(t)$ is injective. This setting encompasses all DMs commonly used in practice, including variance-exploding, variance-preserving (of which the widely used DDPMs [25] are a discretized instance), and sub-variance-preserving [57]. In Appendix A we include explicit formulas for $\psi(t)$, $\sigma^{2}(t)$, and $\lambda(t)$ for these particular DMs.
@@ 3.1 LIDL with a Single Diffusion Model
As opposed to the several normalizing flows used in LIDL which are individually trained on datasets with different levels of noise added, a single DM already works by convolving data with various noise levels and allows density evaluation of the resulting noisy distributions (Equation 5). Hence, we make the observation that LIDL can be used with a single DM. All we need is to relate $\varrho(\cdot,\delta)$ to the density of the DM, $p(\cdot,t)$. In the case of variance-exploding DMs, $\psi(t)=1$, so we can easily use the defining property of the transition kernel in Equation 10 to get
@@
which equals $\varrho(x,\delta)$ from Equation 6 when we choose $t=\sigma^{-1}(e^{\delta})$. [^3] Note that we treat positive and negative superindices differently: e.g. $\sigma^{-1}$ denotes the inverse function of $\sigma$, not $1/\sigma$; on the other hand $\sigma^{2}$ denotes the square of $\sigma$, not $\sigma\circ\sigma$. In turn, we can use LIDL with a single variance-exploding DM by evaluating each $\hat{p}\big{(}x,\sigma^{-1}(e^{\delta_{i}})\big{)}$ through Equation 5. This idea extends beyond variance-exploding DMs; in Appendix B.1 we show that for any arbitrary DM with transition kernel as in Equation 10, it holds that
@@
where $t(\delta)\coloneqq\lambda^{-1}(e^{\delta})$. This equation is relevant because LIDL requires $\log\varrho(\cdot,\delta)$, yet DMs provide $\log p(\cdot,t)$: linking these two quantites as above shows that LIDL can be used with a single DM.
3.2 A Better Implementation of LIDL with a Single Diffusion Model
@@
Using Equation 12 with LIDL still involves computing $\log\hat{p}\big{(}\psi(t(\delta_{i}))x,t(\delta_{i})\big{)}$ through Equation 5 for each $i=1,\dots,m$ before running the regression. Since each of the corresponding ODEs in Equation 4 starts at a different time $t_{0}=t(\delta_{i})$ and is evaluated at a different point $\psi(t(\delta_{i}))x$, this means that a different ODE solver call would have to be used for each $i$, resulting in a prohibitively expensive procedure. To address this, we aim to find an explicit formula for $\partial/\partial\delta\,\log\varrho(x,\delta)$. We do so by leveraging the Fokker-Planck equation associated with Equation 1, which provides an explicit formula for $\partial/\partial t\,p(x,t)$. Using this equation along with the chain rule and Equation 12, we show in Appendix B.2 that, for DMs with transition kernel as in Equation 10,
@@
Then, assuming without loss of generality that $\delta_{1}<\dots<\delta_{m}$, we can use the above equation to define $\log\hat{\varrho}(x,\delta)$ through the ODE
${\textnormal{d}}\log\hat{\varrho}(x,\delta)=\nu\big{(}t(\delta);\hat{s},x\big{)}{\textnormal{d}}\delta,\quad\log\hat{\varrho}(x,\delta_{1})=0.$ (14)
@@
Solving this ODE from $\delta_{1}$ to $\delta_{m}$ produces the trajectory $(\log\hat{\varrho}(x,\delta))_{\delta\in[\delta_{1},\delta_{m}]}$. Since $\nu(t(\delta);\hat{s},x)$ does not depend on $\hat{\varrho}(x,\delta)$, when $\hat{s}=s$ the solution to the ODE above will be off by a constant, i.e., $\log\hat{\varrho}(x,\delta)=\log\varrho(x,\delta)+c_{\text{init}}$ for some $c_{\text{init}}$ that depends on the initial condition of the ODE (00 in this case) but not on $\delta$. Furthermore, while setting the initial condition to 00 might at first appear odd, recall that LIDL fits a regression using $\{(\delta_{i},\log\hat{\varrho}(x,\delta_{i}))\}_{i=1}^{m}$, and thus $c_{\text{init}}$ will be absorbed in the intercept without affecting the slope. In other words, the initial condition is irrelevant, and we can use LIDL with DMs by using a single call to an ODE solver on Equation 14.
3.3 FLIPD: An Efficient Fokker-Planck-Based LID Estimator